A particle of charge q and mass m moves in a circular orbit of radius r with angular speed omega. The ratio of the magnitude of its magnetic moment to that of its angular momentum depents on
  • omega and q
  • omega, q and m
  • q and m
  • omega and m
  •  

Grade:11

1 Answers

Aditya Gupta
4 years ago
As we learnt in 
Magnetic moment (M) -
M=NiA
- wherein
N-number of turns in the coil 
i-current throughout the coil 
A-area of the coil 
Magnetic moment =I.A.=q/T * pir^2= qfpir^2, f is frequency
also, angular momentum =mr^2*w, w is angular freq= 2pif
so, ratio= qfpir^2/mr^2*2pif
q/2m
so it only depends on q and m.
kindly  approve :)